OpenStudy (anonymous):

Solve 2x^2 - 4x = 0 using factoring. Select one: a. {2, 2} Incorrect b. {2, 4} c. {2, -2} d. {0, 2}

OpenStudy (zzr0ck3r):

factor a 2x out and tell me what you have

OpenStudy (science0229):

First, you factor out the 2, so the equation becomes \[2(x ^{2}-2x)=0\]. Now factor the equation inside the parantheses. The common factor is x, so the equation becomes \[2(x(x-2))=0\]. Divide both sides by 2, giving you \[x(x-2)=0\]. Here, either x or x-2 has to be 0 for the equation to be true. So x=0 or x-2=0, which gives you 0 or 2. (0,2). The correct answer is d.